# A Survey on Sequential Recommendation

### v1

@@ 8 Prospects and Future Directions
The SR models have been studied for several years, but there are certain areas that remain to be fully explored.
@@ 8.1 Open-Domain SR
Traditional SR models are trained on a single and static dataset with predefined features. When new items emerge frequently and data resources are increasingly diverse, the traditional SR models are difficult to recommend personalized items. Leveraging the multi-modal features of items and LLMs, open-domain SR models can operate effectively in dynamic environments with a continuous influx of new items. Meanwhile, the open-domain SR models can be transferred across different platforms and domains. By transferring knowledge from external domains or platforms, the open-domain SR models can achieve better recommendation performance within the target domain or platform.
@@ 8.3 Cloud-Edge Collaborative SR
Cloud-edge collaborative SR models combine the strengths of cloud-based and edge-based SR models. Compared with conventional SR models, cloud-edge collaborative SR models can not only improve the recommendation performance but also protect user privacy and leverage resources more effectively. The conventional SR models often require lots of computational resources. Therefore, the conventional SR models are hard to deploy on their own computers. Moreover, transmitting interaction data to a server could risk exposing private information. Therefore, the cloud-edge collaborative SR models are meaningful. Federated SR models are one such approach specifically designed to protect privacy.
@@ 9 Conclusions
In this survey, we have shown that SR models have achieved significant advancements in recommending personalized items for users. From their development, we can have some observations. Firstly, modal-based SR models can tackle the data sparsity and cold-start problems better than pure ID-based SR models. Secondly, modal-based SR models are of high time complexity and need to consume more training time than pure ID-based SR models. Thirdly, modal-based SR models are often associated with better transferability than pure ID-based SR models and SR models combining IDs and general features. Fourthly, multi-modal SR models using IDs can learn fine-grained preferences and achieve more accurate recommendation performance. Lastly, incorporating recent techniques like LLMs into SR models offers further potential for improved recommendation effectiveness. {acknowledgement} We thank the support of National Natural Science Foundation of China (Nos. 62461160311, 62172283 and 62272315), and Guangdong Basic and Applied Basic Research Foundation (Grant No. 2024A1515010122).

### v2

@@ 8 Prospects and Future Directions
SR models have been studied for several years, but there are certain areas that remain to be fully explored.
@@ 8.1 Open-Domain SR
Traditional SR models are trained on a single and static data with predefined features. When new items emerge frequently and data resources are increasingly diverse, the traditional SR models are difficult to recommend personalized items. Leveraging the multi-modal features of items and LLMs, open-domain SR models can operate effectively in dynamic environments with a continuous influx of new items. Meanwhile, the open-domain SR models can be transferred across different platforms and domains. By transferring knowledge from external domains or platforms, the open-domain SR models can achieve better recommendation performance within the target domain or platform.
@@ 8.3 Cloud-Edge Collaborative SR
Cloud-edge collaborative SR models combine the strengths of cloud-based and edge-based SR models. Compared with conventional SR models, cloud-edge collaborative SR models not only can improve the recommendation performance but also protect user privacy and leverage resources more effectively. The conventional SR models often require lots of computational resources. Therefore, the conventional SR models are hard to deploy on their own computers. Moreover, transmitting interaction data to a server could risk exposing private information. Therefore, the cloud-edge collaborative SR models are meaningful. Federated SR models are one such approach specifically designed to protect privacy.
@@ 9 Conclusions
In this survey, we have shown that SR models have achieved significant advancements in recommending personalized items for users. From their development, we can have some observations. Firstly, SR models combining IDs and general features can tackle the data sparsity and cold-start problems better than pure ID-based SR models. Secondly, modal-based SR models are often associated with better transferability than pure ID-based SR models and SR models combining IDs and general features. Thirdly, modal-based SR models are of high time complexity and need to consume more training time than pure ID-based SR models. Fourthly, multi-modal SR models using IDs can learn fine-grained preferences and achieve more accurate recommendation performance. Fifthly, LLMs-based SR models and generative SR models show further potential for improving recommendation effectiveness. Sixthly, SR models can effectively tackle ultra-long interaction sequences by retrieving a few items from these sequences. Finally, data augmentation methods can alleviate the data sparsity and cold-start problems to some extent.
As for future directions, we discuss several interesting topics worthy of exploration, which are about scenario, data, architecture, time, society and explainability. {acknowledgement} We thank the support of National Natural Science Foundation of China (Nos. 62461160311, 62172283 and 62272315), and Guangdong Basic and Applied Basic Research Foundation (Grant No. 2024A1515010122).